# EEG Foundation Models: A Critical Review of Current Progress and Future Directions

### v2

@@ 3.1 Search Strategy
We conducted a comprehensive search across various web platforms, including Google Scholar, arXiv, DBLP, IEEE Xplore, bioRxiv, and medRxiv, to identify relevant research in journals, conferences, workshops, and preprints. We limited our search query to EEG "Foundation Model". We then removed duplicate instances and manually reviewed the title, abstract, and introduction sections to confirm relevant EEG-FMs by identifying phrases similar to "We developed an EEG foundation model…" and "The proposed approach forms the basis for an EEG foundation model…". We note that our search starts from the year 2021 – the year the term Foundation Model was introduced [8] – and includes studies that were published or archived on or before September 30th, 2024.
In addition to the set of foundation models identified above, we also included BrainBERT because of its common utilization as a baseline model in other EEG-FM evaluations [17, 19, 23, 24]. Furthermore, our review does not include foundation models developed on polysomnography (PSG) data [63, 64], which are FMs designed specifically for sleep analysis. While it is possible to extract EEG representations from these FMs, they are not generalizable enough for broader EEG tasks since sleep EEG layouts do not cover the full extent of the head.
@@ 3.2 EEG-FM Summaries
Table 1 summarizes all ten EEG-FMs identified in our search. Although all the EEG-FMs share many commonalities, each FM is unique in its own right. In order to highlight the building blocks of each FM and their unique strengths, below we summarize each FM considering several factors, such as the amount of training data (in channel-hours, calculated as the total recording duration multiplied by the number of EEG channels; see supplement for details), model size (in terms of the number of trainable parameters), the types of EEG data (scalp EEG and/or iEEG), the way inputs are configured (raw time series, power spectra, or time-frequency representation), architectural components (convolutional and/or transformer blocks), the SSL tasks used for pretraining (masked reconstruction, auto-regressive modeling, and/or contrastive learning), and the evaluations performed.
Table 1: Brief model summaries, including training data size, input configurations, data types, architectural components, and SSL tasks. Hyperlinks point to code and model weights, if available.
@@
BrainBERT [15]: As the first released EEG-FM, BrainBERT is relatively smaller than others, with 43.18M parameters and was trained using a modest dataset of 4.5k channel-hours of iEEG data. The inputs were represented as channel-wise spectrograms and a BERT[65]-type model was trained to predict masked patches for different types of spectrograms such as Short-Time Fourier Transform [66](STFT) and Superlets [67]. The model comprised a transformer encoder with a shallow decoder with two linear layers. Evaluations showed generalizability to unseen subjects and unseen electrode locations; however, the test data were from the same distribution as the training data. The evaluations also showed that performing linear probing on BrainBERT embeddings was as good as training supervised deep neural networks (DNN) from scratch for most of the evaluation tasks, which are focused on predicting brain-evoked responses for watching movies. Additionally, their evaluations showed that fine-tuning BrainBERT embeddings can reach DNN performance from scratch with as little as 15% of the training data for one of the tasks. A task-agnostic intrinsic dimensionality [68] (ID)-based analysis showed that the BrainBERT embeddings of different brain regions had different IDs, compared to a relatively constant distribution across electrodes in randomly initialized weights.
@@
Neuro-GPT [16]: This mid-size EEG-FM with 79.53M parameters was trained entirely using scalp EEG data from the full TUH corpus, including 541k channel-hours of clinical scalp EEG data. The model takes raw time series of the 22 EEG channels in the standard 10-20 layout as input and learns EEG representations using a combination of convolution and transformer layers. Those representations are then used as input to a GPT-2 [69] decoder which autoregressively predict the masked latents. It is noteworthy that the decoder has more parameters than the encoder in this setup compared which is not a common practice in other EEG-FMs. This model was evaluated only on EEG data from a BCI motor imagery task with four classes with a different channel configuration than the training data. However, the downstream data were transformed to the original input configuration using an inverse-forward approach[70]. The results showed that fine-tuning or linear probing the pretrained model performed better than models trained from scratch, including some EEG-specific fully-supervised DL approaches.
@@
Brant [17]: Brant is a relatively larger FM with 500M parameters and was trained using 281k channel-hours of iEEG data. However, its pretraining data were limited to a single dataset comprising 9 subjects. This model also takes raw EEG time series as input and is able to take inputs with different channel configurations. The model consisted of a temporal encoder that learns long-term temporal dependencies, a spatial encoder that learns spatial correlations, and a simple linear decoder. The spatial encoder used in this model to capture spatial relationships is a novel contribution compared to previous EEG-FMs. This model also has three scaled-down versions with of 68M, 104M, and 249M parameters, respectively, that are trained on the same data. These models were evaluated on short/long term signal forecasting, frequency phase forecasting, imputation, and seizure detection tasks.
@@
BIOT [18]: This is the smallest of the ten EEG-FMs considered in this review, with only 3.3M parameters, and was pretrained using a contrastive learning objective. The BIOT model introduced a novel approach to take input data with variable lengths and a variable number of channels; it tokenizes each channel into fixed-length segments representing frequency energy vectors, organizes them into “sentences", and uses channel and position embeddings to preserve spatio-temporal information. This model also utilized linear transformers to reduce training time. The model was then evaluated on multiple clinical tasks, such as seizure detection and seizure type classsification, in which the model showed superior performance even without pretraining, and showed even better performance with pretraining, compared to previous fully-supervised DL models.
@@
EEGFormer [19]: This model was also pretrained on the whole TUH corpus, which includes approximately 541k channel-hours of clinical scalp EEG data. It included a transformer encoder and a shallow transformer decoder and it was pretrained using a masked-reconstruction objective. The encoder generated latents of input EEG patches are used to train a vector quantizer to match neural codes generated by a neural codebook. The decoder then reconstructs the original EEG patches using these neural codes. Evaluations included several downstream tasks derived from the TUH corpus and an out-of-distribution (OOD) evaluation on a neonatal seizure detection task. Additional experiments also included interpretability analyses using using the learned codebook.
@@
LaBraM [20]: This model utilized the most diverse, yet a small dataset for pretraining, including 80k channel-hours of scalp EEG data – a subset of the TUH corpus. It was developed in three different scales, with 5.8M, 46M to 369M parameters, respectively. The model consisted of two parts, the neural tokenizer and the neural transformer (LaBraM pretraining model). Both took temporally and spatially patched raw time series data as input and then passed through convolutional temporal encoder of which the outputs are then concatenated with temporal and spatial embeddings and then passed through a set of transformer blocks. The neural tokenizer is trained to reconstruct the amplitude and phase of the patch through a separate decoder, during which the codebook is trained. The neural transformer is trained from scratch, and similar to the neural transformer, except to predict the neural codes inferred using the frozen neural tokenizer, by masked prediction. Although the evaluations demonstrated performance gains various subsets of the TUH corpus, it is unclear whether these results generalize to OOD data.
@@
Mentality [21]: This model aims to capture the complex spatio-temporal dynamics of EEG signals using a Mamba [71]-based state-space model. The architecture of Mentality drew inspiration from other models such as SaShiMi[72], U-Net[73], and EEGNet[74] with the inclusion of Mamba blocks. However, the model was trained and evaluated exclusively on the TUSZ dataset. Furthermore, the unavailability of code or pretrained models limits reproducibility and further evaluations, and makes it inaccessible for broader use as a foundation model.
NeuroLM [22]: This model was inspired by a previous EEG-FM, LaBraM. However, NeuroLM was trained on 7x more data and is one of the largest EEG-FMs with 1.7B parameters, along with 250M and 500M parameter versions. NeuroLM utilized a text-aligned neural tokernizer, which is trained using temporal and frequency reconstruction along with a text/EEG domain classifier that is trained adversarially. The neural tokenizer is similar to that of LaBraM, except for the text alignment component and also included temporal reconstruction in addition to frequency reconstruction. However, despite this novel contribution, the evaluations indicated that the model loses some performance on downstream tasks compared to LaBraM and other state-of-the-art models.
@@
FoME [23]: This model included two versions with 476M and 745M parameters, respectively. The size of the training data was not provided in the manuscript. The model takes masked time series patches and their power spectral densities as input, which are transformed by a temporal encoder. The outputs of the temporal encoder are then reorganized by channels and given as inputs to a spatial encoder to reconstruct masked time patches. FoME was evaluated on multiple downstream tasks, including classification, forecasting, and imputation; however, the evaluations were performed on in-distribution data, limiting any broader conclusions.
@@
BrainWave [24]: This model was pretrained using a large dataset of size 878k channel-hours, including both scalp EEG and iEEG. It includes a transformer encoder and a channel attention module that transform EEG spectrograms into latent representations, which are then decoded by a lightweight decoder. This encoder-decoder architecture was trained using a masked-reconstruction objective. BrainWave is one of the few EEG-FMs trained and evaluated on both scalp EEG and iEEG signals, demonstrating the benefits of joint pretraining over unimodal approaches. The model has been extensively evaluated under different settings, such as cross-subject, cross-hospital, cross-subtype and few-shot classification, showcasing the generalizability and robustness across various clinical tasks.
@@ 4 Comparative Analysis of EEG-FMs
This section surveys the EEG-FMs and compares their design and construction along three major axes covering the pillars of foundation modeling described in Section 2: a) preparation and representation of input data, b) model architecture and self-supervised pretraining, and c) model evaluations. The various considerations along these axes are illustrated in Figure 2. We then highlight the key takeaways from those comparisons.
Figure 2: Overview of comparative analysis. We analyze EEG-FMs along three major axes; input data, modeling and evaluation. EEG data is represented in one of three forms: raw time series, the magnitude power spectrum, and time-frequency representation. The FM architecture may include convolutional blocks to learn low-level patterns and/or transformer blocks to learn higher-level relationships. The FMs are pretrained using SSL approaches; the common SSL approaches used are masked-reconstruction, auto-regressive modeling, and contrastive learning. The pretrained models are then evaluated on various downstream tasks, including clinical and non-clinical tasks. We compare and contrast all ten EEG-FMs along these factors and discuss their strengths and weaknesses.
@@ 4.1 Preparation and Representation of Input Data
Datasets: The data used for pretraining functions as the knowledge base of an FM, making it a highly crucial component in developing a foundation model. Key factors such as diversity, volume, relevance, and, most importantly, quality of the data significantly influence the generalizability of the patterns learned by an FM. Table 2 shows the datasets that were used in the pretraining and evaluation phases of the ten EEG-FMs reviewed in this article.
@@
Most EEG-FMs (e.g., LaBraM, EEGFormer, Neuro-GPT, Mentality, BIOT, and NeuroLM) were trained only on scalp EEG data, although some (e.g., Brant and BrainBERT) were trained exclusively on iEEG data. Two EEG-FMs, FoME and BrainWave, were trained on both scalp and intracranial EEG data. A large portion of the training data in the models utilized scalp EEG was taken from the Temple University Hospital (TUH) EEG corpus [5], which has a total of 541k channel-hours of scalp EEG data. Other scalp EEG datasets used for pretraining include CHB-MIT[75] and various small datasets collected for research purposes. Some FMs utilized EEG data available within PSG datasets such as SHHS[76], Sleep-EDF [77], and CAP Sleep [78]. Additionally, public or private iEEG data accounted for large portions of training data in some models, such as Brant and BrainWave. The common publicly available iEEG datasets used for pretraining were MAYO [79], FNUSA [79], Brain TreeBank[80], and CCEP [81]. Training data volume ranged from 14k channel-hours in BrainBERT to 878k channel-hours in BrainWave. BrainWave utilized the most diverse set of pretraining data, including clinical scalp EEG datasets, sleep EEG datasets, various smaller-scale scalp EEG repositories, and several iEEG datasets. LaBraM and NeuroLM also utilized moderately diverse datasets for pretraining, although all of the datasets were scalp EEG datasets.
@@
Preprocessing and normalization: The content of EEG data is highly sensitive to the choice of pre-processing and normalization steps performed ahead of model training or evaluation. The choices made on resampling, bandpass and notch filtering, artifact removal, and data segmentation into patches or epochs, can directly impact the representations learned by an EEG-FM. Although some models did not comprehensively describe the preprocessing steps, most models applied standard procedures such as bandpass filtering (0.5Hz to a higher cutoff frequency), powerline interference removal via notch filters (50Hz or 60Hz, including harmonics), and downsampling to 200 Hz or 250 Hz to standardize sampling rates across datasets. EEG signals are then segmented into smaller epochs, typically between 1 and 10 seconds. However, none of the EEG-FMs performed explicit artifact removal or exclusion of outlier data, except in cases of expertly-labeled bad data or missing channels. Other less common preprocessing steps included DC offset removal and linear trend removal by FoME and Neuro-GPT. Data normalization practices after preprocessing were not described in several manuscripts. Neuro-GPT applied the commonly used z-transform along the time dimension to normalize the EEG signals, while NeuroLM, BIOT, and LaBraM applied scaling with a constant factor based on the range of input values.
@@
Input representation: EEG is a spatio-temporal data modality, which can be denoted as $X\in\mathbb{R}^{C\times T}$ with $C$ channels and $T$ time points. In addition to the spatial and temporal information, spectral information is also useful in interpreting EEG data, which is used most often as expert-derived features in statistical ML models. Furthermore, time-frequency representations, such as spectrograms or wavelet [82] transforms, are also used to preserve information on all three domains. All EEG-FMs used one or a combination of these representations as input.
@@
The most widely used input representation is the multivariate time series data format, which was adopted by six out of the ten EEG-FMs. With the adoption of transformer architecture, the windows that are segmented from the original recording require further segmentation into smaller patches, also known as tokens. These patches or tokens can then be added with temporal and spatial information with positional embeddings, which allows the models to be trained using EEG data with different channel configurations. This approach was adopted by models such as Brant, NeuroLM, and FoME, which allowed training using datasets with a varying number of channels. Two models, BIOT and EEGFormer, utilized power-spectral data as input instead of the original time series patches, while two other models, Brant and FoME, added power-spectral data to complement the raw time series inputs. Some EEG-FMs, such as LaBraM, Neuro-GPT, and Mentality, standardized different EEG datasets with different channel configurations to a fixed set of input channels. Two other models, BrainBERT and BrainWave, utilized channel-wise time-frequency representations of 1-5 seconds of EEG data as input. Various experiments performed in BrainBERT evaluations utilized spectrograms generated using classical methods like STFT as well as scalograms [82] generated using modern methods such as Superlets.
@@
Table 2: Datasets used to train and evaluate EEG-FMs. We highlight datasets that are not publicly available in bold, clinical and non-clinical datasets in pink and blue backgrounds, respectively, and use superscripts * and # to denote evaluations in internal and external datasets, respectively.
@@
TUEG [83] $train$ $train$ $train$ $train$ $train$
TUAB (Abnormal/normal classification - binary) $eval^{\#}$ $eval^{*}$ $eval^{*}$ $eval^{*}$
TUAR (EEG artifact classification - multiclass) $eval^{*}$ $train$
TUEP (Epilepsy classification - binary) $train$
TUEV (EEG Events classification - multiclass) $eval^{\#}$ $eval^{*}$ $eval^{*}$ $eval^{*}$
TUSL (EEG slowing classification - multiclass) $eval^{*}$ $train$ $eval^{*}$
TUSZ (Seizure type classification - multiclass) $eval^{*}$ $train$ $train/eval^{*}$
CHB-MIT [75] (Peadiatric seizure detection - binary) $eval^{\#}$ $train$ $eval^{\#}$
Sleep-EDFx [77] (Sleep stage classification - multiclass) $train/eval^{*}$ $train$
Siena Scalp EEG [84] (Seizure classification - binary) $train$ $train$ $train$
SHHS [85][76] (Sleep stage classification - multiclass) $train$
PREST (Abnormal event detection - binary) $train$
CAP Sleep [78] (Sleep stage classification - multiclass) $train$
HMC [86] (Sleep stage classification - multiclass) $eval^{\#}$ $train$
SRM [87] (Resting state EEG data) $train$
Schizophrenia-81 (Schizophrenia classification - binary) $train$
Stroke-50 [88] (Hand movement classification - binary) $train$
PD-31 [89] (Parkinson disease classification - binary) $train$
IowaDataset [90] (Unknown task/s) $train$
UNMDataset [91] (Parkinson disease classification - binary) $train$
AD-184 [92] (Alzheimer’s disease classification - binary) $train$
Neaonate dataset [93] (Seizure detection - binary) $eval^{\#}$
IIIC Seizure [94] (IIIC pattern classification) - multiclass $eval^{\#}$

### v3

@@ 3.1 Search Strategy
We conducted a comprehensive search across various web platforms, including Google Scholar, arXiv, DBLP, IEEE Xplore, PubMed, bioRxiv, and medRxiv, to identify relevant research in journals, conferences, workshops, and preprints. We limited our search query, (“EEG” AND “Foundation Model”), to identify FMs that were developed primarily for scalp or intracranial EEG modalities. Our search starts from the year 2021 – the year the term Foundation Model was introduced [bommasani2021opportunities] – and includes studies that were published or archived until September 30th, 2024. We then removed duplicate instances and manually reviewed the title, abstract, and introduction sections to confirm relevant EEG-FMs by identifying phrases to the effect of "We developed an EEG foundation model…" and "The proposed approach forms the basis for an EEG foundation model…". It is important to note that we excluded studies that, a) propose individual components that are building blocks of foundation models (e.g., pretraining strategies, spatio-temporal feature encoders), but not functional off-the-shelf foundation models, and b) finetune existing EEG-FMs to derive task-specific models. Furthermore, our review does not include multi-modal foundation models and sleep foundation models developed using polysomnography (PSG) data [thapa2024sleepfm, ogg2024self]. Nonetheless, we briefly discuss two sleep foundation models published during the same time frame in the supplement for interested readers. Using those search criteria, we identified nine EEG-FMs, Neuro-GPT [cui2024neuro], Brant [zhang2024brant], BIOT [yang2024biot], EEGFormer [chen2024eegformer], LaBraM [jiang_large_2024], Mentality [panchavati2024mentality], NeuroLM [jiang_neurolm_2024], FoME [shi_fome_2024], and BrainWave [yuan2024brainwavebrainsignalfoundation]. In addition, we also included BrainBERT because of its common presence as a baseline in other EEG-FM evaluations [zhang2024brant, chen2024eegformer, shi_fome_2024, yuan2024brainwavebrainsignalfoundation].
@@ 3.2 EEG-FM Summaries
Table 1 summarizes all ten EEG-FMs. Although all the EEG-FMs share many commonalities, each FM is unique in its own right. In order to highlight the building blocks of each FM and their unique strengths, below we summarize each FM considering several factors, such as the amount of training data (in channel-hours, calculated as the total recording duration multiplied by the number of EEG channels; see Supplement Table 1 for details), model size (in terms of the number of trainable parameters), the types of EEG data (scalp EEG and/or iEEG), the way inputs are configured (raw time series, power spectra, or time-frequency representation), architectural components (convolutional and/or transformer blocks), the SSL tasks used for pretraining (masked reconstruction, auto-regressive modeling, and/or contrastive learning), and the evaluations performed.
Table 1: Brief model summaries. We provide brief summaries of the EEG foundation models (EEG-FMs) based on training data size, model size, input configurations, input data type, architectural components, and the self-supervised learning (SSL) tasks used for pretraining. Hyperlinks point to code and model weights, if available. N/A represents cases where channel-hours or model size could not be determined.
@@
BrainBERT [wang2023brainbert]: As the first released EEG-FM, BrainBERT is relatively smaller than others, with 43.18M parameters, and was pretrained using a modest dataset of 4.5k channel-hours of iEEG data. The inputs were represented as channel-wise spectrograms, and a Bidirectional Encoder Representations from Transformers (BERT, [devlin2019bert])-type model was pretrained to predict masked patches for different types of spectrograms based on the Short-Time Fourier Transform (STFT, [gabor1946theory]) and Superlets [moca2021time]. The model comprised a transformer encoder and a shallow decoder with two linear layers. Evaluations focused on evoked brain responses while watching movies [data:wang2024brain] and demonstrated generalizability to unseen subjects and electrode locations; however, the test data were drawn from the same distribution as the training data. The evaluations also showed that linear-probed BrainBERT performance matched that of randomly initialized supervised deep neural networks (DNNs) across most evaluation tasks. Additional evaluations showed that fine-tuned BrainBERT performance could match that of a randomly initialized DNN with as little as 15% of the training data for one task. A task-agnostic intrinsic dimensionality [ansuini2019intrinsic] (ID)-based analysis showed that the BrainBERT embeddings of different brain regions had distinct IDs, whereas the distribution across electrodes was relatively constant under randomly initialized weights.
@@
Neuro-GPT [cui2024neuro]: This mid-size EEG-FM with 79.53M parameters was pretrained entirely using scalp EEG data from the full Temple University Hospital (TUH) EEG Corpus [data:obeid2016temple], which included 541k channel-hours of clinical scalp EEG data. The model takes raw time series of the 22 EEG channels in the standard 10-20 layout as input and learns EEG representations using a combination of convolutional and transformer layers. Those representations are then used as input to a Generative Pretrained Transformer (GPT-2, [radford2019language]) decoder, which autoregressively predicts the masked latents. It is noteworthy that the decoder has more parameters than the encoder in this setup, which is not common in other EEG-FMs. This model was evaluated on EEG data from a 4-class Brain-Computer Interface (BCI) motor imagery task with a different channel configuration than the pretraining data. However, the downstream data were transformed to the original pretraining configuration using an inverse-forward approach [mosher2002eeg]. The results showed that fine-tuning or linear probing the pretrained model performed better than models trained from scratch, including some fully-supervised EEG-DL approaches.
@@
Brant [zhang2024brant]: Brant is a relatively larger FM with 500M parameters and was pretrained using 281k channel-hours of iEEG data. However, its pretraining data were limited to a single dataset comprising 9 subjects. This model accepts raw EEG time series with different channel configurations as input. The model consisted of a temporal encoder that learns long-term temporal dependencies, a spatial encoder that learns spatial correlations, and a simple linear decoder. The spatial encoder used in this model to capture spatial relationships is a novel contribution compared to previous EEG-FMs. This model includes three scaled-down versions with 68M, 104M, and 249M parameters, respectively, all pretrained on the same data. These models were evaluated on short-term and long-term signal forecasting, frequency-phase forecasting, imputation, seizure detection, and pathology detection tasks. The results showed that Brant performs better than baselines in limited-label settings.
@@
BIOT [yang2024biot]: This is the smallest of the ten reviewed EEG-FMs, with only 3.3M parameters, and was pretrained using a contrastive learning objective. The BIOT model introduced a novel approach to handling input data with variable lengths and variable numbers of channels: it tokenizes each channel into fixed-length segments representing frequency vectors, organizes them into ‘sentences’, and uses channel and position embeddings to preserve spatio-temporal information. BIOT used linear transformers to reduce training time and was evaluated across multiple clinical tasks, including seizure detection and seizure type classification. BIOT showed superior performance even without pretraining and even better performance with pretraining, compared to previous supervised EEG-DL models.
@@
EEGFormer [chen2024eegformer]: This model was also pretrained on the full TUH corpus, which includes approximately 541k channel-hours of clinical scalp EEG data. It included a transformer-based encoder and shallow decoder, and was pretrained with a masked reconstruction objective. Latent features of input EEG patches generated by the encoder are used to train a vector quantizer to match neural codes generated by a neural codebook. The decoder then reconstructs the original EEG patches using these neural codes. Evaluations included several downstream tasks derived from the TUH corpus and an out-of-distribution (OOD) evaluation for neonatal seizure detection. Additional experiments included an interpretability analysis using the learned codebook.
@@
LaBraM [jiang_large_2024]: This model used a relatively smaller dataset collected from multiple sources for pretraining, comprising 80k channel-hours of scalp EEG data. It was developed at three parameter scales: 5.8M, 46M, and 369M, respectively. The LaBraM pretrained model consisted of two parts, the neural tokenizer and the neural transformer. The neural tokenizer is first trained to learn a codebook that accurately predicts the amplitude and phase of the raw time series data. Then, with the codebook frozen, the neural transformer is pretrained to predict the codebook-based codes for the masked input patches. Note that this differs from typical pretraining procedures, where the goal is to reconstruct masked patches in the time-domain itself. This transformer model consists of a convolutional temporal encoder and a set of transformer blocks, with temporal and spatial positional encodings used to learn relative relationships between input patches. Although the evaluations demonstrated performance gains in various subsets of the TUH corpus, it is unclear whether these results generalize to out-of-distribution data.
@@
Mentality [panchavati2024mentality]: This model aims to capture the complex spatio-temporal dynamics of EEG signals using a Mamba [gu2023mamba]-based state-space model. The architecture of Mentality drew inspiration from other models such as SaShiMi [goel2022s], U-Net [ronneberger2015u], and EEGNet [lawhern2018eegnet] with the inclusion of Mamba blocks. However, the model was trained and evaluated exclusively on a subset of the TUH corpus.
NeuroLM [jiang_neurolm_2024]: This model was inspired by a previous EEG-FM, LaBraM. However, NeuroLM was trained on 7x more data and is one of the largest EEG-FMs, with 1.7B parameters. Smaller model variants include 250M- and 500M-parameter versions. The neural tokenizer is first trained to learn a codebook that is both aligned with a text embedding space and can accurately reconstruct the input time series data in both time and frequency domains. These codes are then fed into a large language model for multi-channel autoregressive pretraining. Finally, this pretrained model is adapted via joint multi-task instruction tuning, enabling a single model to perform diverse EEG tasks without task-specific fine-tuning. However, despite the novel EEG-text alignment and joint adaptation, NeuroLM performed worse on downstream tasks than LaBraM and other state-of-the-art models.
@@
FoME [shi_fome_2024]: This model included two versions with 476M and 745M parameters, respectively. The size of the training data was not provided in the manuscript. The model takes masked time series patches and their power spectral densities as input, which are then transformed by a temporal encoder. The outputs of the temporal encoder are then reorganized by channel and fed to a spatial encoder to reconstruct masked time patches. FoME was evaluated across multiple downstream tasks, including classification, forecasting, and imputation; however, the evaluations were conducted on in-distribution data, limiting broader conclusions.
@@
BrainWave [yuan2024brainwavebrainsignalfoundation]: This model was pretrained using a large dataset of size 878k channel-hours, including both scalp EEG and iEEG. It includes a transformer encoder and a channel attention module that transform EEG spectrograms into latent representations, which are then decoded by a lightweight decoder. This encoder-decoder architecture was trained using a masked-reconstruction objective. BrainWave is one of the few EEG-FMs trained and evaluated on both scalp EEG and iEEG signals, demonstrating the benefits of joint pretraining over unimodal approaches. The model has been extensively evaluated under different settings, such as cross-subject, cross-hospital, cross-subtype and few-shot classification, showcasing the generalizability and robustness across various clinical tasks.
@@ 4 Comparative Analysis of EEG-FMs
This section surveys the ten EEG-FMs and compares their design and construction along three major axes covering the pillars of foundation modeling described in section 2: a) preparation and representation of input data, b) model architecture and self-supervised pretraining, and c) model evaluations. The various considerations along these axes are illustrated in Figure 2.
Figure 2: Comparative analysis of EEG foundation models (EEG-FMs). Our review analyzes EEG-FMs along three major dimensions; input data configuration, modeling, and evaluation (top figure). A summary of the various approaches undertaken by the EEG-FMs to address those components is shown in the bottom figure. EEG data is represented in one of three forms: raw time series, magnitude power spectrum, and time-frequency representation. Model architecture may include convolutional blocks to learn low-level patterns and/or transformer blocks to learn higher-level relationships. Models are pretrained primarily using self-supervised learning (SSL) approaches; the common SSL approaches used are masked reconstruction, auto-regressive modeling, and contrastive learning. The pretrained models are then evaluated on various downstream tasks, including clinical and non-clinical tasks.
@@ 4.1 Preparation and Representation of Input Data
Datasets: Pretraining data functions as the knowledge base of an FM, making it a highly crucial component. Key factors, such as diversity, data volume, relevance, and, most importantly, data quality, can influence the generalizability of the patterns learned by an EEG-FM. Table 2 shows the datasets that were used in the pretraining and evaluation phases of the ten EEG-FMs reviewed in this article.
@@
Most EEG-FMs (e.g., LaBraM, EEGFormer, Neuro-GPT, Mentality, BIOT, and NeuroLM) were trained only on scalp EEG data, although some (e.g., Brant and BrainBERT) were trained exclusively on iEEG data. Two EEG-FMs, FoME and BrainWave, were trained on both scalp and intracranial EEG data. A large portion of the pretraining data in the models used scalp EEGs from the Temple University Hospital (TUH) EEG corpus [obeid2016temple], which comprises 541k channel-hours of scalp EEG data. This corpus also included smaller derived subsets containing expert annotations for EEG abnormalities (TUAB), seizures (TUSZ), and events (TUEV). Other scalp EEG datasets used for pretraining include CHB-MIT [data:guttag2010chb] and small research datasets. Some FMs utilized EEG data available within PSG datasets such as SHHS [data:zhang2018national], Sleep-EDF [data:kemp2000analysis], and CAP Sleep [data:terzano2001atlas]. Additionally, public or private iEEG data accounted for a large portion of the training data in some models, such as Brant and BrainWave. The common publicly available iEEG datasets used for pretraining were MAYO [nejedly_multicenter_2020], FNUSA [nejedly_multicenter_2020], Brain TreeBank [data:wang2024brain], and CCEP [data:ds004080:1.2.4]. Training data volume ranged from 4.5k channel-hours in BrainBERT to 878k channel-hours in BrainWave. BrainWave utilized the most diverse pretraining dataset, including clinical scalp EEG datasets, sleep EEG datasets, various smaller-scale scalp EEG repositories, and several iEEG datasets. LaBraM and NeuroLM also used moderately diverse pretraining datasets, but were limited to the scalp EEG modality.
@@
Preprocessing and normalization: The content of EEG data is highly sensitive to the preprocessing and normalization steps performed before model training or evaluation. The choices made for resampling, bandpass and notch filtering, artifact removal, and temporal data segmentation into patches or epochs can directly impact the representations learned by an EEG-FM. Although some models did not comprehensively describe the preprocessing steps, most applied standard procedures such as bandpass filtering (0.5Hz to a higher cutoff frequency), powerline interference removal via notch filters (50 Hz or 60 Hz, including harmonics), and downsampling to 200 Hz or 250 Hz to standardize sampling rates across datasets. EEG signals are then segmented into smaller epochs, typically 1-10 seconds long. However, none of the EEG-FMs performed explicit artifact removal or outlier exclusion, except in cases of expertly labeled bad data or missing channels. Other, less common steps included DC offset removal and linear trend removal by FoME and Neuro-GPT. Data normalization practices after preprocessing were not described in several manuscripts. Neuro-GPT applied the commonly used z-transform along the time dimension to normalize EEG signals, whereas NeuroLM, BIOT, and LaBraM used a constant scaling factor based on the input range.
@@
Input representation: EEG is naturally a spatio-temporal data modality recorded using multiple channels. In addition to spatial and temporal information, spectral information is useful for interpreting EEG data, which is most often used as expert-derived features in statistical machine learning (ML) models. Furthermore, time-frequency representations, such as spectrograms and wavelet [grossmann1984decomposition] transforms, are also used to preserve information in all three domains. All EEG-FMs used one or a combination of these representations as input.
@@
The most widely used input representation is the multivariate time-series format, adopted by six of the ten EEG-FMs. With the adoption of a transformer architecture, the windows segmented from the original recording require further segmentation into smaller patches, also known as tokens. These patches or tokens can then be augmented with temporal and spatial information via positional encodings, allowing the models to be trained on EEG data with different channel configurations. This approach was adopted by models such as Brant, NeuroLM, and FoME to pretrain on datasets with varying channel counts. Two models, BIOT and EEGFormer, used power-spectral data as input instead of the original time-series patches, while two other models, Brant and FoME, added power-spectral data to complement the raw time-series inputs. Some EEG-FMs, such as LaBraM, Neuro-GPT, and Mentality, standardized different EEG datasets with different channel configurations to a fixed set of input channels. Two other models, BrainBERT and BrainWave, utilized channel-wise time-frequency representations of 1-5 seconds of EEG data as input. BrainBERT evaluations utilized spectrograms generated using classical methods like STFT, as well as scalograms [grossmann1984decomposition] generated using modern methods such as Superlets.
@@
Table 2: Datasets used for pretraining and evaluation. Here we list the diverse set of datasets and tasks that were used to pretrain and evaluate EEG foundation models (EEG-FMs). We highlight datasets that are not publicly available in bold, clinical and non-clinical datasets in pink and blue backgrounds, respectively, and use superscripts * and # to denote in-distribution and out-of-distribution evaluations, respectively.
@@
TUEG [data:obeid2016temple] traintrain traintrain traintrain traintrain traintrain
TUAB (Abnormal/normal classification - binary) $^{\#} evaleval^{*}$ $^{*} evaleval^{*}$
TUAR (EEG artifact classification - multiclass) $eval^{*}$ traintrain
TUEP (Epilepsy classification - binary) traintrain
TUEV (EEG Events classification - multiclass) $^{\#} evaleval^{*}$ $^{*} evaleval^{*}$
TUSL (EEG slowing classification - multiclass) $eval^{*}$ traintrain $eval^{*}$
TUSZ (Seizure type classification - multiclass) $eval^{*}$ traintrain $train/eval^{*}$
CHB-MIT [data:guttag2010chb] (Peadiatric seizure detection - binary) $eval^{\#}$ traintrain $eval^{\#}$
Sleep-EDFx [data:kemp2000analysis] (Sleep stage classification - multiclass) $eval^{*} traintrain$
Siena Scalp EEG [data:detti2020eeg] (Seizure classification - binary) traintrain traintrain traintrain
SHHS [data:quan1997sleep][data:zhang2018national] (Sleep stage classification - multiclass) traintrain
PREST (Abnormal event detection - binary) traintrain
CAP Sleep [data:terzano2001atlas] (Sleep stage classification - multiclass) traintrain
HMC [data:alvarez2021inter] (Sleep stage classification - multiclass) $eval^{\#}$ traintrain
SRM [data:hatlestad2022bids] (Resting state EEG data) traintrain
Schizophrenia-81 (Schizophrenia classification - binary) traintrain
Stroke-50 [data:liu2024eeg] (Hand movement classification - binary) traintrain
PD-31 [data:ds002778:1.0.5] (Parkinson disease classification - binary) traintrain
IowaDataset [data:anjum2020linear] (Unknown task/s) traintrain
UNMDataset [data:cavanagh2018diminished] (Parkinson disease classification - binary) traintrain
AD-184 [data:vicchietti2023computational] (Alzheimer’s disease classification - binary) traintrain
Neonate dataset [data:stevenson2019dataset] (Seizure detection - binary) $eval^{\#}$
IIIC Seizure [data:ge2021deep] (IIIC pattern classification) - multiclass $eval^{\#}$